OpenStudy (anonymous):

Dividing Radical Expressions Find the value of the square root of 6 over 5 times the square root of 3. in simplest form. 2 over 5. 3 over 4. the square root of 15 over 5. the square root of 2 over 5.

OpenStudy (anonymous):

how did you get that?

OpenStudy (campbell_st):

ok... what about \[\frac{\sqrt{6}}{5\sqrt{3}}\]

OpenStudy (anonymous):

yes that. how do i simplify that

OpenStudy (campbell_st):

rationalise the denominator by multiplying by \[\frac{\sqrt{3}}{\sqrt{3}}\] \[\frac{\sqrt{6}}{5\sqrt{3}}\times \frac{\sqrt{3}}{\sqrt{3}} = \frac{\sqrt{6 \times3}}{5\times \sqrt{3\times3}}\] just simplify and eliminate common factors

OpenStudy (anonymous):

ok so would the final answer be 2 over 3 then?

OpenStudy (anonymous):

i mean 2 over 5

OpenStudy (campbell_st):

no its \[\frac{ 3\sqrt{2}}{ 3 \times5} = \frac {\sqrt{2}}{5}\].

OpenStudy (anonymous):

ok. why does the squared stay there ?

OpenStudy (anonymous):

i dont understand that part

OpenStudy (campbell_st):

find the factors of 18... looking for one that is a square number... \[\sqrt{18} = \sqrt{9\times2} = \sqrt{9} \times \sqrt{2} = 3\sqrt{2}\] 2 is not a square number so needs to remain as a radical

OpenStudy (anonymous):

oooh ok i get it now. thank you.

OpenStudy (campbell_st):

hope it helps